# An Efficient Gloss-Free Sign Language Translation Using Spatial Configurations and Motion Dynamics with LLMs

### v2

@@ Abstract
Eui Jun Hwang Sukmin Cho Junmyeong Lee Jong C. Park School of Computing Korea Advanced Institute of Science and Technology {ehwa20,nelllpic,david516,jongpark}@kaist.ac.kr
@@ 3 Method
We first give an overview of our framework in Section 3.1. We then explain SE and ME in Sections 3.2 and 3.3, respectively. Next, we discuss SA in Section 3.4 and VT-Align in Section 3.5. Finally, we explain the training details in Section 3.6.
@@ 3.1 Framework Overview
Given a sign video $X=\{x_{i}\}_{i=1}^{T}$, where each frame $x_{i}\in\mathbb{R}^{H\times W}$ represents a frame with height $H$ and width $W$, the objective of SLT is to generate a corresponding spoken language sentence $Y=\{y_{j}\}_{j=1}^{U}$, composed of $U$ words. Previous gloss-free methods Zhou et al. (2023); Wong et al. (2024); Gong et al. (2024); Chen et al. (2024) have involved fine-tuning visual encoders using sign language data to make them more domain-specific, leading to improvements in translation accuracy. However, while this fine-tuning introduces more domain knowledge at the feature extraction level, it is unnecessary, especially with LLMs, which already maintain rich visual information from the visual encoder in their latent space Zhang et al. (2024c). Although there may be a trade-off, we argue that utilizing the spatial configurations and motion dynamics with proper alignment and training, offers a more efficient and effective solution.
As shown in Figure 2, Spatial Encoder (SE) and Motion Encoders (ME) extract two distinct features from the sign video $X$: Spatial features $Z_{s}$ capture the spatial configurations Emmorey and Casey (1995), and motion features $Z_{m}$ represent the motion dynamics Bosworth et al. (2019). These features are then integrated into a combined sign feature $Z_{sm}$ via Sign Adapter (SA). The combined feature is then fed to an LLM with a language prompt, guiding the LLM to generate the translation in the desired language. Additionally, we perform Visual-Text Alignment (VT-Align) to minimize the gap between the visual and textual modalities before and during training under SLT supervision. In the following sections, we provide a detailed explanation of SE, ME, SA, and VT-Align.
@@ 3.4 Sign Adapter
In the previous sections, we extracted two distinct visual features: the spatial features $Z_{s}$ and the motion features $Z_{m}$. These features differ in both their dimensions and representation, as depicted in Figure 2. To effectively integrate these features, we introduce an additional module called Sign Adaptor (SA). As shown in Figure 3, SA includes linear projection layers, a 1D CNN, and a Multi-Layer Perceptron (MLP). These components work together to integrate the spatial and motion features into a unified sign representation, denoted as $Z_{sm}$. First, the spatial and motion features are passed through linear projection layers to transform them into features with matching dimensions. Next, the 1D CNN is applied for short-term modeling of the combined features. Finally, a cross-modal MLP Liu et al. (2024a) is employed to bridge the visual and textual modalities. The resulting outputs are represented as $Z_{sm}\in\mathbb{R}^{M\times d^{\prime}}$, where $M$ represents the reduced number of frames after convolution, and $d^{\prime}$ is the dimension aligned with that of the LLM. Although SA aids in bridging the modality gap between visual and textual features during training under the SLT supervision, the gap remains. To tackle this issue, we introduce VT-Align, which will be detailed in the next section.
@@ 4.4 Ablation Study
Next, we explored the impact of different types of LLMs by replacing our model, as shown in Table 4.3. We compared five models, each with a different number of parameters: our method without pre-trained weights, mBART-L, mT0-XL, Flan-T5-XL, and Llama-2. Among these, Flan-T5-XL achieves the best performance, though it nearly ties with mT0-XL. Interestingly, despite its larger parameter count, Llama-2, which is also employed in SignLLM, does not outperform the others. This finding aligns with the observations of Zhang et al., suggesting that scaling up LMs does not always lead to better performance.
@@ Appendix A More Implementation Details
In the SA module, we utilize two distinct linear projection layers tailored for the output feature of ME and SE. For short-term modeling, we employ a 1D CNN configured with a specific sequence of layers: $\{K5,P2,K5,P2\}$, where $K_{\sigma}$ represents a kernel size of $\sigma$, and $P_{\sigma}$ indicates a pooling layer with a kernel size of $\sigma$ Hu et al. (2023). To integrate features into the LLM’s embedding space, we leverage an MLP cross-modal connector Liu et al. (2024a), projecting the features into a 2048-dimensional space.

### v3

@@ Abstract
Eui Jun Hwang Sukmin Cho Junmyeong Lee Jong C. Parkthanks: Corresponding author. School of Computing Korea Advanced Institute of Science and Technology {ehwa20,nelllpic,david516,jongpark}@kaist.ac.kr
@@ 3 Method
We first give an overview of our framework in Section 3.1. We then explain Spatial Encoder (SE) and Motion Encoder (ME) in Sections 3.2 and 3.3, respectively. Next, we discuss Sign Adapter (SA) in Section 3.4 and VT-Align in Section 3.5. Finally, we explain the training details in Section 3.6.
@@ 3.1 Framework Overview
Given a sign video $X=\{x_{i}\}_{i=1}^{T}$, where each frame $x_{i}\in\mathbb{R}^{H\times W}$ represents a frame with height $H$ and width $W$, the objective of SLT is to generate a corresponding translation $Y=\{y_{j}\}_{j=1}^{U}$, composed of $U$ words. Previous gloss-free methods Zhou et al. (2023); Wong et al. (2024); Gong et al. (2024); Chen et al. (2024) have involved fine-tuning visual encoders using sign language data to make them more domain-specific, leading to improvements in translation accuracy. However, while this fine-tuning introduces more domain knowledge at the feature extraction level, it is unnecessary, especially with LLMs, which already maintain rich visual information from the visual encoder in their latent space Zhang et al. (2024c). Although there may be a trade-off, we argue that utilizing the spatial configurations and motion dynamics with proper alignment and training, offers a more efficient and effective solution.
As shown in Figure 2, SE and ME extract two distinct features from the sign video $X$: Spatial features $Z_{s}$ capture the spatial configurations Emmorey and Casey (1995), and motion features $Z_{m}$ represent the motion dynamics Bosworth et al. (2019). These features are then integrated into a combined sign feature $Z_{sm}$ via SA. The combined feature is then fed to an LLM with a language prompt, guiding the LLM to generate the translation in the desired language. Additionally, we perform Visual-Text Alignment (VT-Align) to minimize the gap between the visual and textual modalities before and during training under SLT supervision. In the following sections, we provide a detailed explanation of SE, ME, SA, and VT-Align.
@@ 3.4 Sign Adapter
In the previous sections, we extracted two distinct visual features: the spatial features $Z_{s}$ and the motion features $Z_{m}$. These features differ in both their dimensions and representation, as depicted in Figure 2. To effectively integrate these features, we introduce an additional module called Sign Adaptor (SA). As shown in Figure 3, SA includes linear projection layers, a 1D TCN, and a Multi-Layer Perceptron (MLP). These components work together to integrate the spatial and motion features into a unified sign representation, denoted as $Z_{sm}$. First, the spatial and motion features are passed through linear projection layers to transform them into features with matching dimensions. Next, the 1D TCN is applied for short-term modeling of the combined features. Finally, a cross-modal MLP Liu et al. (2024a) is employed to bridge the visual and textual modalities. The resulting outputs are represented as $Z_{sm}\in\mathbb{R}^{M\times d^{\prime}}$, where $M$ represents the reduced number of frames after convolution, and $d^{\prime}$ is the dimension aligned with that of the LLM. Although SA aids in bridging the modality gap between visual and textual features during training under the SLT supervision, the gap remains. To tackle this issue, we introduce VT-Align, which will be detailed in the next section.
@@ 4.4 Ablation Study
Next, we explored the impact of different types of LLMs by replacing our model, as shown in Table 4.3. We compared five models, each with a different number of parameters: our method without pre-trained weights, mBART-L, mT0-XL, Flan-T5-XL, and Llama-2. Among these, Flan-T5-XL achieves the best performance, though it nearly ties with mT0-XL. Interestingly, despite its larger parameter count, Llama-2, which is also employed in SignLLM, does not outperform the others. This finding aligns with the observations of Zhang et al., suggesting that scaling up LMs does not always lead to better performance. In our case, the main reason likely lies in the fact that larger models generally demand more extensive and higher-quality data to unlock their full potential Kaplan et al. (2020); Hoffmann et al. (2022). Since PHOENIX14T is constrained in both scale and diversity, merely increasing the model size does not necessarily yield substantial performance gains.
@@ Appendix A More Implementation Details
In the SA module, we utilize two distinct linear projection layers tailored for the output feature of ME and SE. For short-term modeling, we employ a 1D TCN configured with a specific sequence of layers: $\{K5,P2,K5,P2\}$, where $K_{\sigma}$ represents a kernel size of $\sigma$, and $P_{\sigma}$ indicates a pooling layer with a kernel size of $\sigma$ Hu et al. (2023). To integrate features into the LLM’s embedding space, we leverage an MLP cross-modal connector Liu et al. (2024a), projecting the features into a 2048-dimensional space.